[Brahms-l] Shift report 20040307 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sun Mar 07 2004 - 07:48:02 EST
------------------------------------------------------------
Date: Sun Mar  7 07:48:01 2004, Shift: 00:00-08:00
Supervisor: EBJ, with: 
Log book#15 pages 78 - 78

Shift Summary:
------------------------------------------------------------
1. We had a lot of beam during the shift.  Right now they are working on storing.

2. We finished up FS at 8deg B Full field.  I had CAS come out to switch the magnet polarities to A for the FS.

3. I didn't encounter any detector problems.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 7 07:48:18 2004

This archive was generated by hypermail 2.1.8 : Sun Mar 07 2004 - 07:48:37 EST